### v1

@@ 5.3.1. Content Organization
The most prominent difference between the two systems is the way content is organized (location-based versus location-agnostic). In the interview, participants confirmed and added that the location was important information about a memory (A3, A4, B1). However, more ambivalent attitudes emerged when it came to the locative interaction mode. Although some acknowledged that it could evoke more memories about the community’s shared context (A10, B1) and encourage more writing (A3, A7, A9, A10) and reading (B4, B5) in daily lives, participants were not always willing to dedicate physical efforts to finding posts in their daily routine, thus seeing fewer posts that are not around their commute routes (A2, A4, A9). The unwillingness to explore had multiple reasons, including personal arrangements (A4, A9), familiarity with the campus (A3), and unattractiveness of the posts (especially at the most commonly used places: canteens and workspaces, see section 5.1) (A2). This is similarly reflected in the questionnaire. The location-based content organization is rated significantly more logical ($U=70.0,p=0.001$, effect size 0.207), navigable ($U=75.0,p=0.001$, effect size 0.213), and linked to the real-world campus environment ($U=110.5,p=0.035$, effect size 0.282); but in the meantime, we could not identify significant differences in people’s impressions ($U=194.0,p=0.696$, effect size 0.532) and motivation ($U=133.5,p=0.156$, effect size 0.344). Concerning the transactive nature of collective memory (wegnerTransactiveMemoryContemporary1987), it highlights the different affordances of the two designs. On the one hand, the locative foregrounds the location-situatedness of the posts (A1, A5, A4, B3) and “‘keeps [things] fresh’ for a longer time” (A4). It may be more engaging if the experience is more tied to the synergy of the content and the physical exertion. On the other hand, the forum design readily delivers the most recent information and provides a more efficient experience (A7, A9, B5, B7). It is more capable of hosting standalone, content-focused narratives.
@@ 5.3.2. UI Style and Interactivity
[c]0.2 Refer to caption    {subcaptionblock}[c]0.6 Refer to caption :
@@ 6.1. Embracing Diverse Interpretations of Collective Memory
Among all the manifold conceptualizations of collection memory, our study especially discovers one debate about the constituents of collective memory that is less focused in previous literature: some participants advocate the retrospective contemplation and emotional expression (which benefit the group with relatable experience and monumental events), whereas some participants emphasize the brief and spontaneous documentation of the current moment as a history for the future (which complete the knowledge of the group’s vibe and easy-to-miss details). This discovery also demonstrates how the HCI perspective and a “bottom-up” approach (section 2.1) adds value to sociological studies of collective memory because it originates from the procedure of communal co-contribution and day-to-day interpersonal interaction. Meanwhile, some previous sociological debates are also present in our findings: one is regarding the community’s monumental or symbolic highlights (schwartzAbrahamLincolnForge2003) versus the minutiae of everyday lives (halbwachsSpaceCollectiveMemory1950); the other is regarding common knowledge and co-experiences (schudsonDynamicsDistortionCollective1995) versus relatable individual experiences and thoughts (zerubavelSocialMindscapesInvitation1997).
@@ 6.2. Interpretative Diversity as a Reflective Design Opportunity
Following our reflection above, we further propose that designers can readily pass the embrace of interpretative diversity on to users. That is, future designs can nudge users to make meaning not only from their preferred styles of narrations and interactions but also from more utterances in the prolific community intelligence. This proposal follows the spirit of reflective design (sengersReflectiveDesign2005): the ultimate goal is not only about researchers’ own reflection on the outcome, but also to continuously bring this concept to users’ conscious awareness and kindle their critical reflection on the constituents of collective memory and their impact.
@@ 6.3. Tensions and Potential Risks of the Design
Collective memory inherently involves the community’s social interactions of memory exchange (wegnerTransactiveMemoryContemporary1987). At a bigger picture, this process is ultimately for the communal good, blending narratives together to form a general impression instead of spotlighting specific individual perspectives (heuxCollectiveMemoryAutobiographical2023; halbwachsSpaceCollectiveMemory1950; harrisCollaborativeRecallCollective2008). However, it is not always the same during individual users’ active co-contribution process. This subsection discusses some potential consequences and risks when designing for collective memory co-contribution. Either hinted by our study participants or drawn from literature, they delineate some important design considerations or interesting research agenda for future HCI scholars.
@@ 6.3.1. Tension between the Communal Cause and the Social Concerns
During the interview, we have observed some participants’ social concerns when engaging in the system (A2, A7): a worry of being judged when a strong social attribute of the platform is perceived (huangSharingFrissonsOnline2024). Such a concern is more salient if the user’s own interpretation of collective memory involves authentic and detailed sharing of personal experiences and feelings. In the scope of collective memory, we note that this tension goes beyond a mere feature choice. It reflects the blurry line between individual content sharing and collective sensemaking, as well as how the social concern may challenge the information gathering stage. As previous “top-down” perspectives often focus on the transactive flow of memories and the formulated collective memory outcome (liuGrassrootsHeritageCrisis2010; wegnerTransactiveMemoryContemporary1987; halbwachsSpaceCollectiveMemory1950; lakCollectiveMemoryUrban2019), our grassroots findings complement the HCI bigger picture with the stage before memories are exchanged and collectively processed. In this regard, we also call for more future design explorations to lower the mental barrier—for example, making use of gamification (liuCoArgueFosteringLurkers2023), accentuating the aggregated summarization instead of individuals’ utterances (huangSharingFrissonsOnline2024; robinsonChatHasNo2022), channeling more social acknowledgement and reciprocity (caraban23WaysNudge2019).
@@ 6.3.2. Potential Privacy Concerns of Geolocation Data
Apart from the social concerns, the richer feature set in Drop’n’Collect also leads to more possibilities of privacy concerns. For example, the geo-tagging of personal memories and photos may inherently encode sensitive information, such as identifiers, location and activity histories, and absence information, potentially resulting in unconscious information leaks (ruizvicenteLocationRelatedPrivacyGeoSocial2011; freniPreservingLocationAbsence2010). Such concerns are not present in our interview even though we asked the participants explicitly, possibly because we have addressed them during the onboarding briefing. Still, future real-world deployments are suggested to clearly state the potential risks of involving sensitive geolocation data, ensuring users’ privacy awareness (barthPrivacyParadoxInvestigating2017). In addition, these deployments can also provide granular control over the information visibility and anonymity, satisfying varying privacy preferences of different users in different communities. For example, the system can offer to configure the users to whom a memory post is visible, mask faces in photographs when viewed by different people, and replace original voices with AI-generated ones.
@@ 6.3.3. Potential Negative Social Phenomena
In our study, we have set a clear rule that toxic, hostile, and other negative social acts in both systems are strictly prohibited. However, in real-world deployments, social factors are inherently attributed to such collective memory facilitation systems (as also addressed in section 6.3.1), and negative social acts may inevitably occur, either consciously or unconsciously. For example, people may have conflicting perspectives and memories regarding the same event (houYouDontUnderstand2025; momennejadBridgeTiesBind2019), and one person’s memory may be another person’s trauma (zhangCrisisCollectiveMemory2020). Besides, within an enlarged user base, in-groups and out-groups may form, resulting in memory content barely understood outside subgroups (maattaEverydayDiscourseSpace2021), or an increased toxicity level (parkYouChangeWay2024). Thus, if a community platform is dedicated to collective memory development, it is suggested to actively detect potential harmful and toxic content before publishing, and prompt the user to refine it.
@@ 6.4. Locative Interactions for Community Members’ Devoted Narration
The first point is based on the significantly more user activities at canteens and workspaces because of participants’ “resting” state at those locations. Therefore, such physical affordances of different locations can be an effective element to utilize in future locative systems. It is similar to Millard et al.’s utilization of the terrain’s geological traits (millardCanyonsDeltasPlains2013)—except that they focused on content consumption experiences and did not consider the functional affordances of geolocations. In our case, for example, designers can specially enhance the “resting” locations with extra motivation and authoring preparation—as if checkpoints in games.
Secondly, considering the physical efforts of initiating one locative narration session, authoring several posts in a dedicated session may be more preferable than fragmented writing in everyday routine. Dedicated authoring sessions can be functionally supported by the ability to create more fluent experiences, such as “memory trajectories”. They can serve as both a logical thread connecting narrative pieces and the building block for immersive exploratory experiences (cheverstSupportingConsumptionCoauthoring2017; lovlieYouAreOne2012; benfordInteractionTrajectoriesDesigning2009). Furthermore, designers can build up users’ intrinsic motivation through reward systems like gamification or amplified social feedback (liuCoArgueFosteringLurkers2023), as well as an enhanced recommendation system based on users’ narration activities and writing styles.
Several existing HCI works have shown the power of locative UGC platforms for the community’s cause, as the content readily augments the geographical reality and takes part in placemaking (cranshawJourneysNotesDesigning2016; cheverstSupportingConsumptionCoauthoring2017). But many either focused on short and ad-hoc content (juBreadcrumbSNSAsynchronous2015; bowserPrototypingPLACEScalable2013; mcgookinStudyingDigitalGraffiti2014; changTourgetherExploringTourists2020) or did not evaluate the experience of serious content authoring (cranshawJourneysNotesDesigning2016; cheverstSupportingConsumptionCoauthoring2017). Therefore, our design process and findings also generally add insights to the mobile HCI community on effectively supporting locative and devoted narrations.
@@ Acknowledgements.
This project is partially supported by the Hong Kong SAR Research Grants Council’s Theme-based Research Grant Scheme (Project No. T43-518/24N).

### v2

@@ 5.3.1. Content Organization
The most prominent difference between the two systems is the way content is organized (location-based versus location-agnostic). In the interview, participants confirmed and added that the location was important information about a memory (A3, A4, B1). However, more ambivalent attitudes emerged when it came to the locative interaction mode. Although some acknowledged that it could evoke more memories about the community’s shared context (A10, B1) and encourage more writing (A3, A7, A9, A10) and reading (B4, B5) in daily lives, participants were not always willing to dedicate physical efforts to finding posts in their daily routine, thus seeing fewer posts that are not around their commute routes (A2, A4, A9). The unwillingness to explore had multiple reasons, including personal arrangements (A4, A9), familiarity with the campus (A3), and unattractiveness of the posts (especially at the most commonly used places: canteens and workspaces, see section 5.1) (A2). This is similarly reflected in the questionnaire. The location-based content organization is rated significantly more logical ($U=70.0,p=0.001$, effect size 0.207), navigable ($U=75.0,p=0.001$, effect size 0.213), and linked to the real-world campus environment ($U=110.5,p=0.035$, effect size 0.282); but in the meantime, we could not identify significant differences in people’s impressions ($U=194.0,p=0.696$, effect size 0.532) and motivation ($U=133.5,p=0.156$, effect size 0.344). Concerning the transactive nature of collective memory (Wegner, 1987), it highlights the different affordances of the two designs. On the one hand, the locative foregrounds the location-situatedness of the posts (A1, A5, A4, B3) and “‘keeps [things] fresh’ for a longer time” (A4). It may be more engaging if the experience is more tied to the synergy of the content and the physical exertion. On the other hand, the forum design readily delivers the most recent information and provides a more efficient experience (A7, A9, B5, B7). It is more capable of hosting standalone, content-focused narratives.
@@ 5.3.2. UI Style and Interactivity
[c]0.2 Refer to caption    {subcaptionblock}[c]0.6 Refer to caption A memory postcard about the sunset and the backs of a cuddling couple. The two check-in photographs are of users making funny, teasing body gestures at the couples in the postcard.
@@ 6.1. Embracing Diverse Interpretations of Collective Memory
Among all the manifold conceptualizations of collection memory, our study especially discovers one debate about the constituents of collective memory that is less focused in previous literature: some participants advocate the retrospective contemplation and emotional expression (which benefit the group with relatable experience and monumental events), whereas some participants emphasize the brief and spontaneous documentation of the current moment as a history for the future (which complete the knowledge of the group’s vibe and easy-to-miss details). This discovery also demonstrates how the HCI perspective and a “bottom-up” approach (section 2.1) adds value to sociological studies of collective memory because it originates from the procedure of communal co-contribution and day-to-day interpersonal interaction. Meanwhile, some previous sociological debates are also present in our findings: one is regarding the community’s monumental or symbolic highlights (Schwartz, 2003) versus the minutiae of everyday lives (Halbwachs, 1950); the other is regarding common knowledge and co-experiences (Schudson, 1995) versus relatable individual experiences and thoughts (Zerubavel, 1997).
@@ 6.2. Interpretative Diversity as a Reflective Design Opportunity
Following our reflection above, we further propose that designers can readily pass the embrace of interpretative diversity on to users. That is, future designs can nudge users to make meaning not only from their preferred styles of narrations and interactions but also from more utterances in the prolific community intelligence. This proposal follows the spirit of reflective design (Sengers et al., 2005): the ultimate goal is not only about researchers’ own reflection on the outcome, but also to continuously bring this concept to users’ conscious awareness and kindle their critical reflection on the constituents of collective memory and their impact.
@@ 6.3. Tensions and Potential Risks of the Design
Collective memory inherently involves the community’s social interactions of memory exchange (Wegner, 1987). At a bigger picture, this process is ultimately for the communal good, blending narratives together to form a general impression instead of spotlighting specific individual perspectives (Heux et al., 2023; Halbwachs, 1950; Harris et al., 2008). However, it is not always the same during individual users’ active co-contribution process. This subsection discusses some potential consequences and risks when designing for collective memory co-contribution. Either hinted by our study participants or drawn from literature, they delineate some important design considerations or interesting research agenda for future HCI scholars.
@@ 6.3.1. Tension between the Communal Cause and the Social Concerns
During the interview, we have observed some participants’ social concerns when engaging in the system (A2, A7): a worry of being judged when a strong social attribute of the platform is perceived (Huang et al., 2024). Such a concern is more salient if the user’s own interpretation of collective memory involves authentic and detailed sharing of personal experiences and feelings. In the scope of collective memory, we note that this tension goes beyond a mere feature choice. It reflects the blurry line between individual content sharing and collective sensemaking, as well as how the social concern may challenge the information gathering stage. As previous “top-down” perspectives often focus on the transactive flow of memories and the formulated collective memory outcome (Liu, 2010; Wegner, 1987; Halbwachs, 1950; Lak and Hakimian, 2019), our grassroots findings complement the HCI bigger picture with the stage before memories are exchanged and collectively processed. In this regard, we also call for more future design explorations to lower the mental barrier—for example, making use of gamification (Liu et al., 2023), accentuating the aggregated summarization instead of individuals’ utterances (Huang et al., 2024; Robinson et al., 2022), channeling more social acknowledgement and reciprocity (Caraban et al., 2019).
@@ 6.3.2. Potential Privacy Concerns of Geolocation Data
Apart from the social concerns, the richer feature set in Drop’n’Collect also leads to more possibilities of privacy concerns. For example, the geo-tagging of personal memories and photos may inherently encode sensitive information, such as identifiers, location and activity histories, and absence information, potentially resulting in unconscious information leaks (Ruiz Vicente et al., 2011; Freni et al., 2010). Such concerns are not present in our interview even though we asked the participants explicitly, possibly because we have addressed them during the onboarding briefing. Still, future real-world deployments are suggested to clearly state the potential risks of involving sensitive geolocation data, ensuring users’ privacy awareness (Barth and De Jong, 2017). In addition, these deployments can also provide granular control over the information visibility and anonymity, satisfying varying privacy preferences of different users in different communities. For example, the system can offer to configure the users to whom a memory post is visible, mask faces in photographs when viewed by different people, and replace original voices with AI-generated ones.
@@ 6.3.3. Potential Negative Social Phenomena
In our study, we have set a clear rule that toxic, hostile, and other negative social acts in both systems are strictly prohibited. However, in real-world deployments, social factors are inherently attributed to such collective memory facilitation systems (as also addressed in section 6.3.1), and negative social acts may inevitably occur, either consciously or unconsciously. For example, people may have conflicting perspectives and memories regarding the same event (Hou et al., 2025; Momennejad et al., 2019), and one person’s memory may be another person’s trauma (Zhang et al., 2020). Besides, within an enlarged user base, in-groups and out-groups may form, resulting in memory content barely understood outside subgroups (Määttä et al., 2021), or an increased toxicity level (Park et al., 2024). Thus, if a community platform is dedicated to collective memory development, it is suggested to actively detect potential harmful and toxic content before publishing, and prompt the user to refine it.
@@ 6.4. Locative Interactions for Community Members’ Devoted Narration
The first point is based on the significantly more user activities at canteens and workspaces because of participants’ “resting” state at those locations. Therefore, such physical affordances of different locations can be an effective element to utilize in future locative systems. It is similar to Millard et al.’s utilization of the terrain’s geological traits (Millard et al., 2013)—except that they focused on content consumption experiences and did not consider the functional affordances of geolocations. In our case, for example, designers can specially enhance the “resting” locations with extra motivation and authoring preparation—as if checkpoints in games.
Secondly, considering the physical efforts of initiating one locative narration session, authoring several posts in a dedicated session may be more preferable than fragmented writing in everyday routine. Dedicated authoring sessions can be functionally supported by the ability to create more fluent experiences, such as “memory trajectories”. They can serve as both a logical thread connecting narrative pieces and the building block for immersive exploratory experiences (Cheverst et al., 2017; Løvlie, 2012; Benford et al., 2009). Furthermore, designers can build up users’ intrinsic motivation through reward systems like gamification or amplified social feedback (Liu et al., 2023), as well as an enhanced recommendation system based on users’ narration activities and writing styles.
Several existing HCI works have shown the power of locative UGC platforms for the community’s cause, as the content readily augments the geographical reality and takes part in placemaking (Cranshaw et al., 2016; Cheverst et al., 2017). But many either focused on short and ad-hoc content (Ju et al., 2015; Bowser et al., 2013; McGookin et al., 2014; Chang et al., 2020) or did not evaluate the experience of serious content authoring (Cranshaw et al., 2016; Cheverst et al., 2017). Therefore, our design process and findings also generally add insights to the mobile HCI community on effectively supporting locative and devoted narrations.
@@ Acknowledgements.
This project is partially supported by the Hong Kong SAR Research Grants Council’s Theme-based Research Grant Scheme (Project No. T43-518/24-N).